OpenStudy (anonymous):

A little help: 7m − 5 − 13m = 25 I got down to -6m - 5= 25 I know you take the 5 and put it on the other side of the 25 but I cant remember if you add the 5 or subtract it

OpenStudy (phi):

you don't *really* move the 5 with an equation (to keep it equal) you can *do the same thing to both sides* For example we could add + 5 to both sides, like this -6m - 5 + 5= 25+5 now simplify You could (but don't want to) add -5 to both sides: -6m -5 - 5 = 25 - 5 but that is not helpful, because you get -6m - 10 = 20 (not what you want)

OpenStudy (anonymous):

Okay thank you sooo much! You made this a lot easier! I need help on another, in a couple of minutes do you think you could help?

OpenStudy (phi):

please make it a new post.